# Bar properties as a function of wavelength in Illustris TNG50: analysis of mock images

### v1

@@ Abstract
Recent studies used Illustris TNG50 galaxies to study bar formation, evolution, and properties like length, strength, and pattern speed. In simulations, these are typically derived from particle positions and mass distribution, neglecting stellar light and extinction effects. However, observational studies indicate that bar appearance depends on wavelength. To test whether this dependence exists in TNG50 at $z\approx 0$, we analysed 50 strongly-barred galaxies using mock images from SKIRT radiative transfer simulations covering infrared to ultraviolet filters (Spitzer $3.6\,\upmu\text{m}$, SDSS i, r, g, S-PLUS J0378, GALEX NUV, and GALEX FUV). Bar ellipticity and length were measured via ellipse fitting. Ellipticity generally increases by 6 per cent from $3.6\,\upmu\text{m}$ to $g$ band, and by 9 per cent to $J0378$ band. On average, TNG50 bars cannot be said to be longer in bluer filters when the entire sample is used. However, the trend is indeed detected when only star-forming galaxies are considered. In this star-forming subsample, bar length increases by 10 per cent from $3.6~{}\mu\text{m}$ to $g$ band, and by 17 per cent to $J0378$ band; moreover, the bar can appear up to 20–30 per cent longer in bluer mocks than in the mass map. Over 90 per cent of bars vanish in UV due to minimal emission by dominant stellar populations. We reproduced the bar properties morphological dependence phenomenon using age-filtered mass maps, where older stars form shorter, rounder bars, and younger stars generate longer, more elliptical ones. TNG50 bars exhibit a wavelength-dependent trend similar to observations: bars more elliptical and longer in bluer filters, with this effect being stronger in star-forming galaxies.
@@
pubyear: 2025pagerange: Bar properties as a function of wavelength in Illustris TNG50: analysis of mock images27
@@ 1 Introduction
A recent observational paper by Menéndez-Delmestre et al. (2024) investigates the behaviour of measuring galactic bars in the local universe across different observational filters, ranging from infrared to ultraviolet (Galex NUV, Galex FUV, Johnson B, Johnson R, Spitzer IRAC/$3.6\,\upmu\text{m}$). The aim of Menéndez-Delmestre et al. (2024) is to understand the effect of wavelength-dependent characteristics of the same bar to extrapolate findings to band-shifting scenarios at high redshift. The methodology is based on measuring bar properties by fitting ellipses (Jedrzejewski, 1987), where the maximum ellipticity determines bar length. The main result that the authors present is the wavelength dependence of bar length and strength (characterized by ellipticity) with observed band: they find that in bluer filters bands tend to show bars that are more elongated and elliptical (by around 10 per cent), while the same bar appears shorter and less elliptical in infrared. Hypotheses to explain this dependency have been proposed, including the possible influence of star-forming clumps at the bar ends or the effect of kinematic fractionation (Athanassoula et al., 2017; Fragkoudi et al., 2017; Debattista et al., 2017; Neumann et al., 2020), which results in younger stars being preferentially located along the length of the bar, leading to an elongation in bluer filters. Additionally, the contribution of a bulge in redder filters could reduce bar ellipticity, affecting its isophotes (Menéndez-Delmestre et al., 2024; Díaz-García et al., 2016).
@@
Cosmological simulations play an important role in studying the formation and evolution of the fraction of barred galaxies. The properties and evolution of bars have been explored using data from simulations such as EAGLE (Algorry et al., 2017; Cavanagh et al., 2022) and IllustrisTNG, in both TNG100 (Zhao et al., 2020) and TNG50 (Rosas-Guevara et al., 2022). Notably, when comparing the bar fractions from simulations with observational results, there is a good agreement at redshift $z=0$. However, this fraction tends to diverge at higher redshift: while simulations like IllustrisTNG show a bar fraction close to 40 per cent (Rosas-Guevara et al., 2022) at redshift $z=3$, recent observational results from JWST indicate a significant decrease in this fraction (Guo et al., 2024).
@@
To provide context, we can refer to works that employ mock images based on cosmological simulations. Baes et al. (2024a) created an atlas of TNG50 mock galaxies at various wavelengths, each with different orientations relative to the cosmological cube. This atlas, generated using SKIRT, features noise-free simulations. An example analysis included evaluating the effective radius of galaxies across different wavelengths (Baes et al., 2024b). Another example of creating mock images of individual galaxies is presented in Rodriguez-Gomez et al. (2019), which uses Illustris simulations processed with SKIRT (Camps & Baes, 2020) to model observations from Pan-STARRS (Survey Telescope and Rapid Response System; Chambers et al. 2016). These simulations include steps for adding synthetic noise, enhancing their resemblance to real observational data. A distinct application is presented in Snyder et al. (2023), which uses galaxies from IllustrisTNG to simulate light cone mosaics analogous to those observed with JWST. By concatenating multiple galaxies into a single field of view, this approach models deep field observations across a series of broadband filters utilized by JWST-NIRCam and the Hubble Space Telescope cameras.
@@
Among large cosmological simulations, TNG50 shows a bar fraction at $z=0$ that is most consistent with observations. In contrast, the original Illustris (Peschken & Łokas, 2019) and Eagle (Algorry et al., 2017) simulations report bar fractions of around 20 per cent at $z=0$, and New Horizon (Reddish et al., 2022) shows no barred galaxies at this redshift. The Auriga simulation (Fragkoudi et al., 2025) exhibits comparably good results for the bar fraction at $z=0$ and a better representation of bar fraction evolution. Due to its zoom-in nature, it focuses on 40 Milky Way-like halos, while TNG50 evaluates the bar fraction in hundreds of halos that formed spontaneously within a broader cosmological context.
@@ 2 Simulations and methods
This section presents the sample of barred galaxies in the IllustrisTNG cosmological simulation, as well as the methods used to process images of the mocks created with SKIRT and measure the morphological properties of the bars.
@@ 2.1 The IllustrisTNG Simulation
The IllustrisTNG project (Pillepich et al., 2018b; Nelson et al., 2019) provides a suite of simulations that represent the current state-of-the-art of galaxy formation and evolution (Pillepich et al., 2018a; Weinberger et al., 2017) based on hydrodynamical cosmological models. It is performed using the moving-mesh code AREPO (Springel, 2010) considering a flat $\Lambda$CDM cosmological model, containing additional physics algorithms in addition to the gravo-magneto-hydrodynamical code such as star formation, colling, feedback from supernovae and active galactic nuclei. The IllustrisTNG simulation executed runs within three cubic cosmological volumes, each having side lengths of 300, 100, and 50 cMpc, respectively, with each volume featuring progressively finer spatial resolutions. This setup allows not only the formation of large cosmological structures, but also the study of subhalos containing galaxies in detail, enabling morphological analyses of their components. Examples of such analyses can be found in the works of Tacchella et al. (2019) and Zana et al. (2022).
@@ 2.2 Generating mock images
The radiative transfer (RT) simulation stage, which considers the stars of the TNG50 galaxies as the source and simulates the dust properties generated from the gas in the subhalos, was performed using the Monte Carlo RT code SKIRT (Baes et al., 2011; Camps & Baes, 2020). SKIRT calculates physical processes, including scattering, absorption, and emission through the transfer medium (Baes & Camps, 2015), and features an import interface compatible with data formats from hydrodynamical simulations like IllustrisTNG. This allows the import of stellar information as radiative sources and the derivation of the spatial distribution and properties of dust from the gas particles. In recent years, SKIRT has been consistently used for RT calculations with dust absorption, using particles from the IllustrisTNG simulation as input (e.g. Rodriguez-Gomez et al., 2019; Baes et al., 2024b, a; Popping et al., 2022).
@@
Using stellar particle data and dust distribution derived from the gas component of the TNG50 galaxies, data cubes were generated with bands covering the entire wavelength range from 0.1 to 5 $\mu$m from the RT simulations. The images generated by SKIRT are saved in FITS format (Flexible Image Transport System). By using the Python toolkit for working with SKIRT (PTS) library (Camps & Baes, 2020), it is possible to generate images in specific observational bands by applying appropriate transmission curves. For our analysis, we generated images in the in the infrared using the Spitzer IRAC/$3.6\,\upmu\text{m}$ channel, along with SDSS i, r, g filters, and the SPLUS narrowband J0378 (near SDSS u). We also explored ultraviolet filters, using Galex FUV and Galex NUV (see Appendix 19). The standard SDSS u band proved unsuitable for our analysis due to the clumpy appearance of its mocks from star-forming regions, which severely affected ellipse fitting. Similar issues were encountered with the GALEX FUV and NUV filters discussed in Subsection 4.3. Therefore, the SPLUS J0378 narrowband filter was used as a viable alternative to the problematic SDSS u band, providing a shorter wavelength coverage than SDSS g but without effects of excessive star-forming clumps.
@@ 2.4 Ellipse fitting and bar properties
Various methods are employed to measure the length and ellipticity of the bar, both in observations and in simulated galaxies. Among the most commonly used methods are Fourier analysis, as exemplified in the study by Rosas-Guevara et al. (2022) on IllustrisTNG galaxies, and the fitting of ellipses to isodensity curves(Jedrzejewski, 1987; Wozniak et al., 1995; Erwin, 2005), as utilized in Zhao et al. (2020) and Lu et al. (2024), also with IllustrisTNG data. Fourier analysis is effective in measuring bars by decomposing the light or projected mass distribution into symmetric components, while ellipse fitting is widely used to characterize the shape of the bar based on the spatial distribution of isodensities, providing a detailed measure of ellipticity and length. A comparative analysis of these methods has been explored in observational and numerical contexts, as discussed in Michel-Dansac & Wozniak (2006), highlighting their strengths and limitations in different regimes.
@@ 3 Results
In this section, we analyse the wavelength dependence of bar properties in our simulated galaxies. We examine how bar ellipticity and length vary across different mock observation filters, identifying key trends and significant dependencies within specific galaxy subsamples. We also explore the implications for bar detectability in the UV mocks and correlation between these morphological dependencies and stellar age distributions.
@@ 3.1 Comparison of wavebands across individual galaxies
We applied the bar measurement method to all 50 galaxies in our sample, considering the J0378, g, r, i bands, the Spitzer IRAC/$3.6\,\upmu\text{m}$, and the mass maps. For each galaxy, we obtained a comparison profile of the ellipticity across the different filters, which defines both the bar length and its maximum ellipticity. An example of this result is shown in Fig. 4, where we present an example of the ellipse fit on the isophotes of a barred galaxy (subhalo ID566365), analysing the ellipticity profiles in the 2D mass maps and mocks generated using the five observational filters. The solid lines in Fig. 4 show the ellipticity variations of the fitted isophotes as a function of the semi-major axis for each filter, with the dashed lines indicating the estimated bar length in each band. The bar lengths vary between 3.25 kpc for the mass, 3.95 kpc for Spitzer $3.6\,\upmu\text{m}$, 4.10 kpc for i-, 4.20 kpc for r-, 4.40 kpc for g- and 5.05 kpc for J038-band.
Comparison between bar length measurements using the Fourier decomposition method from Rosas-Guevara et al. (2022) and the ellipse fitting method (this work) can be found in Fig. 5. The distribution of points reveals that, for short bars, the Fourier method tends to produce values both above and below those obtained via ellipse fitting. However, for longer bars ($\geq 4$ kpc), the Fourier method systematically yields larger measurements, this is evident in both the mass maps and when compared to observational filters. Linear fits (represented by dashed lines in the same colour as the data points in Fig. 5) emphasizes this trend, on average, all filters show longer bar lengths when measured with the ellipse fitting method compared to Fourier analysis. This offset is most prominent in the bluer filters (J0378 and g-band), while redder filters and the mass maps yield bar length measurements that are more consistent among themselves. A detailed explanation of the methodology used to compare the bar length measurements across filters is provided in Sections 3.1 and 3.2. The physical mechanisms associated with the host galaxy that help explain these systematic differences are discussed in Section 3.3.
@@
Figure 5: Comparison between bar length measurements using the Fourier decomposition method from Rosas-Guevara et al. (2022) and the ellipse fitting method (this work). The purple, blue, green, red, orange and black symbols correspond to the bar length measured in the J0378-, g-, r-, i-bands, the IRAC/$3.6\,\upmu\text{m}$ and mass maps respectively. The solid line represents the one-to-one relation. Dashed lines of matching colour indicate linear fits to the data points in each band.
@@ 3.2 Monotonicity test between wavelengths
To assess whether the differences in measurements identified in the previous section exhibit any monotonic dependence between the analysed filters, we undertake a series of tests. The tests used to evaluate monotonicity and the slope of this trend are Spearman’s Rank Correlation Test and a linear slope fitting test. In other words, after establishing that bars are generally longer and more elliptical in a sequential trend with wavelength, we seek to quantify how systematic this dependence is.
The use of Spearman’s test in this context involves comparing the bar ellipticity and length measurements by fixing the sequence of observational filters along the $x$-axis in order of increasing wavelength. Measurements derived from the 2D mass maps were not included in the Spearman’s test. We begin with the Spitzer IRAC/$3.6\,\upmu\text{m}$ followed by SDSS i, r, g and ending with the shortest wavelength filter, SPLUS J0378. This arrangement allows us to evaluate the potential relationship between the measurements and the variation in wavelength.
Our initial hypothesis is that the ellipticity measurements will show a monotonic increase as we progress to filters with shorter wavelengths (from the $3.6\,\upmu\text{m}$ to the observational band J0378). However, as illustrated in Fig. 10, where we display the ellipticities measured in the $3.6\,\upmu\text{m}$, i-, r-, g- and J0378-bands for 3 galaxies in our sample, the trend with band can vary. When the trend is strictly postiive monotonic, the Spearman coefficient will be equal to 1.00, indicating a perfect correlation. In cases where the monotonicity is weaker or absent, the coefficient will be smaller than 1.00, reflecting a quasi-monotonic or non-monotonic trend. For the purpose of Spearman’s test, the filter names are treated as dimensionless integers along the $x$-axis. It is important to note that equivalent examples to these described for a strong positive trend also exist for a strong negative monotonic trend. A Spearman coefficient greater than 0.8 (or less than $-0.8$ for a negative trend) is typically considered to indicate a strong monotonic trend.
@@ 3.3 Correlations with the host galaxy
As an illustration of this correlation, we plotted the slope of ellipticity as a function of the specific star formation rate of the host galaxies (Fig. 14). This analysis was performed for both the entire sample of 50 galaxies and a subset consisting only of galaxies with a Spearman coefficient greater than 0.8, indicating a strong monotonic dependence on wavelength. In both cases, the correlation suggests a clear trend: galaxies with higher star formation activity tend to exhibit a stronger wavelength-dependent effect, while halos with lower star formation rates are typically the ones showing non-monotonic behaviour and lower slopes. The phenomenon studied appears to manifest more notably in star-forming galaxies and is less present in passive galaxies.
Figure 14: Correlation between specific star formation rate and the slope of bar ellipticity measured across bands. Each point represents one of the 50 galaxies in our sample. Galaxies with a strong positive monotonic trend in ellipticity are highlighted in blue. Galaxies with a strong negative monotonic trend in ellipticity are highlighted in red. The remaining galaxies are shown in gray.
Fig. 15 presents a two-panel comparison of the fitted slope of bar-length variation across bands against two key parameters, revealing bimodality. In the upper panel, the slope is plotted versus the bar length measured in the $J0378$ filter ($L_{J0378}$): galaxies with shorter $L_{J0378}$ (< 4 kpc) predominantly exhibit negative monotonicity, meaning their bars appear systematically larger in redder filters, whereas those with longer $L_{J0378}$ show positive monotonicity, with bars systematically larger in bluer filters. This behaviour arises because the $J0378$ filter preferentially traces star-forming regions in the bar, so intrinsically longer bars in this filter, hosting more star formation, display a stronger wavelength-dependent lengthening. In the lower panel, the same slope is compared to sSFR, again separating the sample into two subgroups: galaxies with $\log(\mathrm{sSFR}/\mathrm{yr}^{-1})\geq-11$ preferentially exhibit negative monotonicity, while those with $\log(\mathrm{sSFR}/\mathrm{yr}^{-1})<-11$ more often positive monotonicity. This split highlights two distinct galaxy populations, defined by their star-formation activity, that exhibit opposite wavelength-dependent bar-length trends.
Figure 15: Two-panel comparison of the fitted slope of bar-length measurements across bands for our 50-galaxy sample. Top panel: slope versus bar length measured in the J0378 filter ($L_{J0378}$). Bottom panel: slope versus specific star formation rate ($\log\,\mathrm{sSFR}/\mathrm{yr}^{-1}$). In both panels, blue points denote galaxies with a strong positive monotonic trend, red points denote those with a strong negative trend, and gray points show the remainder.

### v2

# Bar properties as a function of wavelength in TNG50: analysis of mock images

@@ Abstract
Recent studies used TNG50 galaxies to study bar formation, evolution, and properties like length, strength, and pattern speed. In simulations, these are typically derived from particle positions and mass distribution, neglecting stellar light and extinction effects. However, observational studies indicate that bar appearance depends on wavelength. To test whether this dependence exists in TNG50 at $z\approx 0$, we analysed 50 strongly-barred galaxies using mock images from SKIRT radiative transfer simulations covering infrared to ultraviolet filters (Spitzer $3.6\,\upmu\text{m}$, SDSS i, r, g, S-PLUS J0378, GALEX NUV, and GALEX FUV). Bar ellipticity and length were measured via ellipse fitting. Ellipticity generally increases by 6 per cent from $3.6\,\upmu\text{m}$ to $g$ band, and by 9 per cent to $J0378$ band. On average, TNG50 bars cannot be said to be longer in bluer filters when the entire sample is used. However, the trend is indeed detected when only star-forming galaxies are considered. In this star-forming subsample, bar length increases by 10 per cent from $3.6~\mu\text{m}$ to $g$ band, and by 17 per cent to $J0378$ band; moreover, the bar can appear up to 20–30 per cent longer in bluer mocks than in the mass map. Over 90 per cent of bars vanish in UV due to minimal emission by dominant stellar populations. We reproduced the bar properties morphological dependence phenomenon using age-filtered mass maps, where older stars form shorter, rounder bars, and younger stars generate longer, more elliptical ones. TNG50 bars exhibit a wavelength-dependent trend similar to observations: bars more elliptical and longer in bluer filters, with this effect being stronger in star-forming galaxies.
@@
pubyear: 2025pagerange: Bar properties as a function of wavelength in TNG50: analysis of mock images27
@@ 1 Introduction
[…]
@@
Cosmological simulations play an important role in studying the formation and evolution of the fraction of barred galaxies. The properties and evolution of bars have been explored using data from simulations such as EAGLE (Algorry et al., 2017; Cavanagh et al., 2022) and IllustrisTNG, in both TNG100 (Zhao et al., 2020) and TNG50 (Rosas-Guevara et al., 2022). Notably, when comparing the bar fractions from simulations with observational results, there is a good agreement at redshift $z=0$. However, this fraction tends to diverge at higher redshift: while simulations like TNG50 show a bar fraction close to 40 per cent (Rosas-Guevara et al., 2022) at redshift $z=3$, recent observational results from JWST indicate a significant decrease in this fraction (Guo et al., 2024).
@@
To provide context, we can refer to works that employ mock images based on cosmological simulations. Baes et al. (2024a) created an atlas of TNG50 mock galaxies at various wavelengths, each with different orientations relative to the cosmological cube. This atlas, generated using SKIRT, features noise-free simulations. An example analysis included evaluating the effective radius of galaxies across different wavelengths (Baes et al., 2024b). Another example of creating mock images of individual galaxies is presented in Rodriguez-Gomez et al. (2019), which uses TNG simulations processed with SKIRT (Camps & Baes, 2020) to model observations from Pan-STARRS (Survey Telescope and Rapid Response System; Chambers et al. 2016). These simulations include steps for adding synthetic noise, enhancing their resemblance to real observational data. A distinct application is presented in Snyder et al. (2023), which uses galaxies from TNG to simulate light cone mosaics analogous to those observed with JWST. By concatenating multiple galaxies into a single field of view, this approach models deep field observations across a series of broadband filters utilized by JWST-NIRCam and the Hubble Space Telescope cameras.
@@
Among large cosmological simulations, TNG50 and TNG100 shows a bar fraction at $z=0$ that is most consistent with observations. This improvement over its predecessor, Illustris, is largely attributed to updates in the subgrid physics model in the IllustrisTNG suite, particularly in the treatment of AGN and stellar feedback (Zhou et al., 2020). In contrast, the original Illustris (Peschken & Łokas, 2019) and Eagle (Algorry et al., 2017) simulations report bar fractions of around 20 per cent at $z=0$, and New Horizon (Reddish et al., 2022) shows no barred galaxies at this redshift. The Auriga simulation (Fragkoudi et al., 2025) exhibits comparably good results for the bar fraction at $z=0$ and a better representation of bar fraction evolution. Due to its zoom-in nature, it focuses on 40 Milky Way-like halos, while TNG50 evaluates the bar fraction in hundreds of halos that formed spontaneously within a broader cosmological environment.
@@ 2 Simulations and methods
This section presents the sample of barred galaxies in the TNG50 cosmological simulation, as well as the methods used to process images of the mocks created with SKIRT and measure the morphological properties of the bars.
@@ 2.1 The IllustrisTNG Simulation
[…]
@@ 2.2 Generating mock images
The radiative transfer (RT) simulation stage, which considers the stars of the TNG50 galaxies as the source and simulates the dust properties generated from the gas in the subhalos, was performed using the Monte Carlo RT code SKIRT (Baes et al., 2011; Camps & Baes, 2020). SKIRT calculates physical processes, including scattering, absorption, and emission through the transfer medium (Baes & Camps, 2015), and features an import interface compatible with data formats from hydrodynamical simulations like IllustrisTNG. This allows the import of stellar information as radiative sources and the derivation of the spatial distribution and properties of dust from the gas particles. In recent years, SKIRT has been consistently used for RT calculations with dust absorption, using particles from the TNG50 simulation as input (e.g. Rodriguez-Gomez et al., 2019; Baes et al., 2024b, a; Popping et al., 2022).
@@
[…]
@@ 2.4 Ellipse fitting and bar properties
Various methods are employed to measure the length and ellipticity of the bar, both in observations and in simulated galaxies. Among the most commonly used methods are Fourier analysis, as exemplified in the study by Rosas-Guevara et al. (2022) on TNG50 galaxies, and the fitting of ellipses to isodensity curves(Jedrzejewski, 1987; Wozniak et al., 1995; Erwin, 2005), as utilized in Zhao et al. (2020) and Lu et al. (2024), also with IllustrisTNG data. Fourier analysis is effective in measuring bars by decomposing the light or projected mass distribution into symmetric components, while ellipse fitting is widely used to characterize the shape of the bar based on the spatial distribution of isodensities, providing a detailed measure of ellipticity and length. A comparative analysis of these methods has been explored in observational and numerical contexts, as discussed in Michel-Dansac & Wozniak (2006), highlighting their strengths and limitations in different regimes.
@@ 3 Results
In this section, we analyse the wavelength dependence of bar properties in our simulated galaxies. We examine how bar ellipticity and length vary across different mock observation filters, identifying key trends and significant dependencies within specific galaxy subsamples. We also explore the implications for bar detectability in the UV mocks and correlation between these morphological dependencies and stellar age distributions. A detailed explanation of the methodology used to compare the bar length measurements across filters is provided in Sections 3.1 and 3.2. The physical mechanisms associated with the host galaxy that help explain these systematic differences are discussed in Section 3.3.
@@ 3.1 Comparison of wavebands across individual galaxies
[…]
Comparison between bar length measurements using the Fourier decomposition method from Rosas-Guevara et al. (2022) and the ellipse fitting method (this work) can be found in Fig. 5. The distribution of points reveals that, for short bars, the Fourier method tends to produce values both above and below those obtained via ellipse fitting. However, for longer bars ($\geq 4$ kpc), the Fourier method systematically yields smaller measurements, this is evident in both the mass maps and when compared to observational filters. Linear fits (represented by dashed lines in the same colour as the data points in Fig. 5) emphasizes this trend, on average, all filters show longer bar lengths when measured with the ellipse fitting method compared to Fourier analysis. This offset is most prominent in the bluer filters (J0378 and g-band), while redder filters and the mass maps yield bar length measurements that are more consistent among themselves.
@@
[…]
@@ 3.2 Monotonicity test between wavelengths
To assess whether the differences in measurements identified in the previous section exhibit any monotonic dependence between the analysed filters, we undertake a series of tests. The tests used to evaluate monotonicity and the slope of this trend are Spearman’s Rank Correlation Test and a linear slope fitting test. In other words, we aim to quantify whether the variations in bar length and ellipticity across filters follow a systematic trend with wavelength.
[…]
Our initial hypothesis is that the ellipticity measurements will show a monotonic increase as we progress to filters with shorter wavelengths (from the $3.6\,\upmu\text{m}$ to the observational band J0378). However, as illustrated in Fig. 10, where we display the ellipticities measured in the $3.6\,\upmu\text{m}$, i-, r-, g- and J0378-bands for 3 galaxies in our sample, the trend with band can vary. When the trend is strictly postiive monotonic, the Spearman coefficient will be equal to 1.00, indicating a perfect correlation. In cases where the monotonicity is weaker or absent, the coefficient will be smaller than 1.00, reflecting a quasi-monotonic or non-monotonic trend. For the purpose of Spearman’s test, the filter names are treated as dimensionless integers along the $x$-axis. It is important to note that equivalent examples to these described for a strong positive trend also exist for a strong negative monotonic trend. A Spearman coefficient greater than 0.8 (or less than 0.8-0.8 for a negative trend) is typically considered to indicate a strong monotonic trend.
@@ 3.3 Correlations with the host galaxy
As an illustration of this correlation, we plotted the slope of ellipticity as a function of the specific star formation rate of the host galaxies (Fig. 14). This analysis was performed for both the entire sample of 50 galaxies and a subset consisting only of galaxies with a Spearman coefficient greater than 0.8, indicating a strong monotonic dependence on wavelength. In both cases, the correlation suggests a clear trend: galaxies with higher star formation activity tend to exhibit a stronger wavelength-dependent effect, while galaxies with lower star formation rates are typically the ones showing non-monotonic behaviour and lower slopes. The phenomenon studied appears to manifest more notably in star-forming galaxies and is less present in passive galaxies.
Figure 14: Correlation between specific star formation rate and the slope of bar ellipticity measured across bands. Each point represents one of the 50 galaxies in our sample. Galaxies with a strong positive monotonic trend in ellipticity are highlighted in blue. Galaxies with a strong negative monotonic trend in ellipticity are highlighted in red. The remaining galaxies are shown in gray. The Spearman correlation shown refers to the full sample.
Fig. 15 presents a two-panel comparison of the fitted slope of bar-length variation across bands against two key parameters, revealing bimodality. In the upper panel, the slope is plotted versus the bar length measured in the $J0378$ filter ($L_{J0378}$): galaxies with shorter $L_{J0378}$ (< 4 kpc) predominantly exhibit negative monotonicity, meaning their bars appear systematically larger in redder filters, whereas those with longer $L_{J0378}$ show positive monotonicity, with bars systematically larger in bluer filters. This behaviour arises because the $J0378$ filter preferentially traces star-forming regions in the bar, so intrinsically longer bars in this filter, hosting more star formation, display a stronger wavelength-dependent lengthening. In the lower panel, the same slope is compared to sSFR, again separating the sample into two subgroups: galaxies with $\log(\mathrm{sSFR}/\mathrm{yr}^{-1})<-11$ preferentially exhibit negative monotonicity ($n=20$), while those with $\log(\mathrm{sSFR}/\mathrm{yr}^{-1})\geq-11$ more often positive monotonicity ($n=18$). This split highlights two distinct galaxy populations, defined by their star-formation activity, that exhibit opposite wavelength-dependent bar-length trends. The Spearman correlation coefficient shown in each panel refers to the full 50-galaxy sample.
Figure 15: Two-panel comparison of the fitted slope of bar-length measurements across bands for our 50-galaxy sample. Top panel: slope versus bar length measured in the J0378 filter ($L_{J0378}$). Bottom panel: slope versus specific star formation rate ($\log\,\mathrm{sSFR}/\mathrm{yr}^{-1}$). In both panels, blue points denote galaxies with a strong positive monotonic trend, red points denote those with a strong negative trend, and gray points show the remainder. The Spearman correlation shown refers to the full sample.